OpenStudy (anonymous):

What number would be missing? Please help. x*28.5*102=4355.4

OpenStudy (anonymous):

solve for x

OpenStudy (anonymous):

approx 1.5

OpenStudy (anonymous):

I would need an exact number though.

OpenStudy (anonymous):

its pretty long...

OpenStudy (anonymous):

Well not exact. Lol. Somewhere near precise I guess you could say.

OpenStudy (anonymous):

1.4982456140350877192982456140351 how's that?

OpenStudy (anonymous):

if it's meters then 2 decimal places should be enough

OpenStudy (anonymous):

x*28.5*102=4355.4 \[x = \frac{4355.4}{28.5 \times 102} = \frac{4355.4}{2907} = 1.498 \approx 1.5\] @Krazygio

OpenStudy (anonymous):

Thank you!